OpenStudy (anonymous):

18x + 92 = 281 - 11x ?

OpenStudy (anonymous):

get your x's and constants on different sides

OpenStudy (curry):

189/29 = x

OpenStudy (anonymous):

18x+92=281-11x Since -11x contains the variable to solve for, move it to the left-hand side of the equation by adding 11x to both sides. 18x+92+11x=281 Since 18x and 11x are like terms, add 11x to 18x to get 29x. 29x+92=281 Since 92 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 92 from both sides. 29x=-92+281 Add 281 to -92 to get 189. 29x=189 Divide each term in the equation by 29. (29x)/(29)=(189)/(29) Simplify the left-hand side of the equation by canceling the common terms. x=(189)/(29)

OpenStudy (anonymous):

x = 189/29

OpenStudy (aroub):

11x+18x=-92+281 29x=189 x=189/29

OpenStudy (anonymous):

-281 + 92= -18x -11x -189 = -29x x= 189/29

OpenStudy (anonymous):

it's saying my answer options are (A) 3 (B) 6 and 15/29 (C)18 and 5/9 or (D) 27 How do I get any of those out of 189/29?

OpenStudy (anonymous):

the answer is B) since 6 + 15/29 = 189/29

OpenStudy (anonymous):

ohhh. I see what you mean.

OpenStudy (anonymous):

just divide 189 by 29

OpenStudy (anonymous):

29 times 6 is 174 189-174=15 so 6 15/29